in the formula, V= kq/r, what is k, q and r . Pls explain where and why is this formula usd and show the derivation too!!!!!! Pls!!!!!

Grade:9

2 Answers

vikas askiitian expert
13 years ago

V = potential at any point ...

V = kq/r 

another definition =>  workdone in bringing +ve test charge(q=1)  from infinity to a point which is at a distance r , near a charge q ...

F = k(q)(qo)/r2                      (from culambs law)

qo is +ve test charge , F is force bw test charge & q charge...

F = -dU/dr

dU/dr = -F = -kqqo/r2

dU = -kqqo(1/r2)dr             lim from infinity to r

U = -kqqo(-1/r)                 lim infinity to r                     [   (1/r2)dr = -1/r     ]

U = kqqo/r                  (qo = 1)

so , U = kq/r

here K is columbs constant , k = 9*109units , q is charge ...

this is the required result...

vikas askiitian expert
13 years ago

approve if u like my ans